# Junyu Cao and Mohsen Bayati

### v1

@@ 3 One-dimensional Human Comparison
Once the low-dimensional representation $\hat{\varphi}$ has been established in the initial stage, as detailed in Section 4, we introduce a framework for model alignment through human comparisons. The parameter $\boldsymbol{\theta}$ may lie in a multi-dimensional space. To illustrate our algorithm, we first describe the one-dimensional human comparison procedure, and then extend to the multi-dimensional scenario.
The primary challenge in designing the framework lies in strategically selecting pairwise comparisons to minimize sample complexity, which bears resemblance to the principles of active learning. In this section, we first discuss the deterministic selection (with 100% accuracy) and then study the probabilistic selection. For the probabilistic selection, we propose an algorithm named Alignment through Probabilistic Bisection (APB) and characterize its sample complexity.
@@ 3.1 Deterministic Bisection
The overall goal is to use the least samples to find some $\theta$ such that $d(\theta,\theta^{*})\leq\varepsilon$. When human is able to select a better model with 100% accuracy, the deterministic bisection algorithm can be utilized to approach the true parameter. The algorithm operates by maintaining a search interval interval $[\theta^{-},\theta^{+}]$ that contains the true parameter $\theta^{*}$. The interval is initialized by $[-\beta_{\Theta},\beta_{\Theta}]$, and at round $k$ the center of this interval with respect to the distance $d(\cdot,\cdot)$ is used as query point. Specifically, defining the query point $\theta_{k}$ to be the single point in the set $H(\theta^{-},\theta^{+})$, we ask the human to compare utility of two points, denoted by $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$, that have equal distance $\Delta/2$ from $\theta_{k}$ in opposite directions, where $\Delta$ is called comparison granularity and is a tuning parameter. For example, for the Euclidean distance, at each round $\theta_{k}=(\theta^{-}+\theta^{+})/2$, and $c_{\Delta}^{-}(\theta_{k})=\theta_{k}-\Delta/2$ and $c_{\Delta}^{+}(\theta_{k})=\theta_{k}+\Delta/2$. If the human selects $c_{\Delta}^{-}(\theta_{k})$ then we eliminate the upper half of the search interval by updating $\theta^{+}$ to be equal to $\theta_{k}$; otherwise, we eliminate the bottom half of the search interval by setting $\theta^{-}$ to be equal to $\theta_{k}$. After this bisection step, we continue the process by selecting $\theta_{k+1}$ to be $H(\theta^{-},\theta^{+})$ again, and repeat the above till the length of the search interval is below the precision $\varepsilon$. The pseudo code for this is presented in Algorithm 1.
@@
Figure 1: An illustrative example. Among two choices $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$, $c_{\Delta}^{-}(\theta_{k})$ would be selected since it is closer to the true parameter. In this case, the interval $(\theta_{k},\theta^{+}]$ is eliminated.
@@
Suppose human makes the selection with 100% accuracy $(\gamma\to 0)$. Fix $\varepsilon>0$. After $O(\log_{2}\left(\frac{\beta_{\Theta}}{\varepsilon}\right))$ rounds, Algorithm 1 reaches a point $\theta_{k}$ such that $d(\theta_{k},\theta^{*})\leq\varepsilon$.
However, the deterministic model of human selection described above exhibits two significant limitations: firstly, in practical scenarios, humans may not always be able to identify the correct choice, implying that $\gamma>0$; secondly, the accuracy of human selections is influenced by the comparison granularity $\Delta$. To address these concerns, we introduce a more general probabilistic model in the following subsection.
@@ 3.2 Probabilistic Bisection
First, we enrich the model to capture the probabilistic nature of human comparisons by noting that, in the RUM, when $\gamma$ is strictly positive, human’s selection is random with certain distribution that favors the correct answer, parameterized by $\gamma$. Precisely, Lemma 2.6 indicates that human will select $c_{\Delta}^{-}(\theta)$ with probability
$p^{-}_{\gamma}(\theta):=\left[1+\exp\left\{\frac{u(c_{\Delta}^{+}(\theta))-u(c_{\Delta}^{-}(\theta))}{\gamma}\right\}\right]^{-1}\,,$
and selects $c_{\Delta}^{+}(\theta)$ with probability $p^{+}_{\gamma}(\theta)$ equal to $1-p^{-}_{\gamma}(\theta)$.
We note that, in the above equation, the selection probability depends on the comparison granularity. This dependence, naturally captures difficulty of the selection for the human. Specifically, when $\Delta$ is too small, $c_{\Delta}^{-}(\theta)$ and $c_{\Delta}^{+}(\theta)$ converge to each other, hence $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ converge to $1/2$, representing the difficulty of distinguishing between too highly similar choices. On the other hand, when $\Delta$ is too large, one expects the selection to become difficult as well as both options are far away from $\theta^{*}$. This can be captured for example when the distance function converges to a finite value for far away points from $\theta^{*}$ as in Example 2.4. In this case, $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ could also converge to 1/2 (see Example 3.2). In the remaining of the paper, for brevity, we do not include $\Delta$ in the notations $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$.
@@
Consider $u(\theta)=-\frac{1}{1+\exp(-|\theta|)}$ with the highest utility achieved at $\theta=0$. When $\Delta$ is too large, both $u(c_{\Delta}^{+}(\theta))$ and $u(c_{\Delta}^{-}(\theta))$ are close to -1 even $\theta$ is far away from 0, and thus $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ are both close to 1/2.
@@
Let $\varepsilon>0$ denote the precision level and $\delta>0$ represent the confidence level. The $(\varepsilon,\delta)$-alignment problem aims to identify a $\theta$ such that $\mathbb{P}(d(\theta,\theta^{*})\leq\varepsilon)\geq 1-\delta$.
@@
Our objective is to solve the $(\varepsilon,\delta)$-alignment problem with small sample complexity. The algorithm design is inspired by the probabilistic bisection developed in Frazier et al. (2019). While their work primarily focuses on the analysis of convergence rates, our emphasis is on conducting a sample complexity analysis for an algorithm designed to terminate when both the precision level and confidence level criteria are met.
@@
If $\theta<\theta^{*}$, then $p_{\gamma}^{-}(\theta)<1/2$ and $p_{\gamma}^{+}(\theta)>1/2$; vice versa. We introduce some notation for simplicity. Let $Y(\theta)$ denote the selection: $Y(\theta)=c_{\Delta}^{-}(\theta)$ with probability $p_{\gamma}^{-}(\theta)$ and $Y(\theta)=c_{\Delta}^{+}(\theta)$ with probability $p_{\gamma}^{+}(\theta)$. Define $\tilde{Z}(\theta)=2\cdot\boldsymbol{1}(Y(\theta)=c_{\Delta}^{+}(\theta))-1$, so that $\tilde{Z}(\theta)=1$ if $Y(\theta)=c_{\Delta}^{+}(\theta)$ (so that we believe $\theta^{*}$ more likely to be to the right of $\theta$) and $\tilde{Z}(\theta)=-1$ if $Y(\theta)=c_{\Delta}^{-}(\theta)$ (so that we believe $\theta^{*}$ more likely to be to the left of $\theta$). Note that we have $\tilde{Z}(\theta)=1$ with probability $p_{\gamma}^{+}(\theta)$. Also, define
@@
$\tilde{p}(\theta)=\begin{cases}\mathbb{P}(\tilde{Z}(\theta)=1),&\mbox{if }\theta\leq\theta^{*}\\ \mathbb{P}(\tilde{Z}(\theta)=-1),&\mbox{if }\theta>\theta^{*}\end{cases}\,,$
so that $\tilde{p}(\theta)$ is the probability that $\tilde{Z}(\theta)$ correctly indicates the direction of the root $\theta^{*}$ from $\theta$.
@@
Figure 2 shows the structure of introducing our main algorithm. The general idea of the Algorithm APB is described as follows. There are two types of moves: the horizontal moves and the vertical moves. For any fixed $\theta$, we ask human to repeatedly evaluate $c_{\Delta}^{-}(\theta)$ and $c_{\Delta}^{+}(\theta)$. We refer this as the vertical moves. Once the vertical stopping criteria is satisfied, the algorithm decides the next $\theta$ to evaluate according to the updated distribution. We call this step as the horizontal moves. The horizontal moves continue until the horizontal stopping criteria is satisfied. In the following two sections, we discuss horizontal moves and vertical steps, as well as the stopping criteria, in sequence. Then we introduce the two-stage Human-AI collaboration framework for multiple dimensions in Section 4 and practical implementations in Section 5.
@@
One-dimensional APB (Section 3.2) Horizontal moves (Section 3.2.2); Complexity characterization (Section 3.3) Sample complexity (Theorem 3.5) Two-stage framework for Human-AI collaboration (Section 4.1); complexity characterization (Section 4.3) Practical implementations (Section 5) Vertical moves (Section 3.2.1); Complexity characterization (Section 3.4)
Figure 2: Roadmap for introducing 2.
@@ 3.2.1 Vertical moves
Fix some $\theta$. Repeatedly, we ask human evaluations of the query point $\theta$, achieved through comparisons of $c^{+}_{\Delta}(\theta)$ and $c^{-}_{\Delta}(\theta)$. We observe a sequence of signs $\{\tilde{Z}_{i}(\theta)\}_{i}$. If $\theta<\theta^{*}$, then the expectation $\mathbb{E}[\tilde{Z}_{i}(\theta)]=2\tilde{p}(\theta)-1>0$; if $\theta>\theta^{*}$, then the expectation $\mathbb{E}[\tilde{Z}_{i}(\theta)]=1-2\tilde{p}(\theta)<0$. From the response $\tilde{Z}_{i}(\theta)$, we define the simple random walk $S(\theta)=\{S_{m}(\theta):m\geq 0\}$ with $S_{0}(\theta)=0$ and $S_{m}(\theta)=\sum_{i=1}^{m}\tilde{Z}_{i}(\theta)$ for $m\geq 1$. Sequential test of power one indicates whether the drift $\omega$ of a random walk satisfies the hypothesis $\omega<0$ versus $\omega>0$.
Define the stopping time $\tau^{\uparrow}_{1}(\theta)=\inf\{m\in\mathbb{N}:|S_{m}(\theta)|\geq\hbar_{m}\}$ where $\{\hbar_{m}\}_{m}$ is a positive sequence. Here $\uparrow$” ($\rightarrow$) represents “vertical” (horizontal, respectively). The test decides that $\theta<\theta^{*}$ if $S_{\tau_{1}^{\uparrow}(\theta)}(\theta)\geq\hbar_{\tau^{\uparrow}_{1}(\theta)}$, that $\theta>\theta^{*}$ if $S_{\tau_{1}^{\uparrow}(\theta)}(\theta)\leq-\hbar_{\tau^{\uparrow}_{1}(\theta)}$ and the test does not make a decision if $\tau^{\uparrow}_{1}(\theta)=\infty$.
Assume now that the algorithm queries at some point $\theta_{k}\neq\theta^{*}$ at the $(k+1)^{th}$ iteration. We then observe a random walk with $m^{th}$ term $S_{k,m}=S_{k,m}(\theta_{k})=\sum_{i=1}^{m}Z_{k,i}(\theta_{k})$ until we reach the end of the power-one test. Define the new signal
@@
Define $\hbar_{m}=(2m(\ln(m+1)-\ln\eta))^{1/2}$. On the event $\theta_{k}>\theta^{*}$, $\mathbb{P}(Z_{k}(\theta_{k})=+1|\theta_{k})\leq\eta/2$. On the event $\theta_{k}<\theta^{*}$, $\mathbb{P}(Z_{k}(\theta_{k})=-1|\theta_{k})\leq\eta/2$. That is, the detection accuracy $p(\theta)$ when the vertical move stops is bounded by $p(\theta)\geq 1-\eta/2$.
Lemma 3.4 states that the detection accuracy is lower bounded by $p_{c}:=1-\eta/2$ where the accuracy can be chosen through the selection of $\eta$. As $\theta$ approaches $\theta^{*}$, the differentiation between choices becomes increasingly challenging because the drift tends to diminish, potentially requiring more steps in the testing process. However, it is crucial to remember that our objective is to locate a $\theta$ such that $|\theta-\theta^{*}|\leq\varepsilon$ with a confidence level of $1-\delta$. In the vicinity of $\theta^{*}$, waiting until $|S_{m}(\theta)|\geq\hbar_{m}$ to terminate the test may not be necessary. A high volume of queries at a particular point $\theta_{k}$ suggests close proximity between $\theta_{k}$ and $\theta$, making it advisable to consider terminating the test sooner.
Based on this insight, we introduce the stopping time $\tau^{\uparrow}(\delta,\varepsilon)$ that is determined by the precision and confidence level. This stopping time regulates the maximum number of queries carried out at a single point as $\theta$ approaches the local neighborhood of the true parameter (a concept to be elaborated on in Section 3.4). Specifically, as $\theta$ converges towards $\theta^{*}$ within a defined local neighborhood, denoted as $\{\theta^{\prime}:d(\theta,\theta^{\prime})\leq a\}$ for some positive constant $a$, surpassing the threshold $\tau^{\uparrow}(\delta,\varepsilon)$ signifies the difficulty in distinguishing between the parameters due to their proximity. The algorithm terminates.

### v2

@@ 4 One-dimensional Human Comparison
The low-dimensional representation $\hat{\varphi}$ is established in the initial stage, as we discuss in Section 5. Following this stage, our framework calls for model alignment through human comparisons. At this point, the parameter $\boldsymbol{\theta}$ may lie in a multi-dimensional space. To illustrate our algorithm, we first describe the one-dimensional human comparison procedure and then extend it to the multi-dimensional scenario.
The primary challenge in designing the framework lies in strategically selecting pairwise comparisons to minimize the number of comparisons (or sample complexity), which resembles the principles of active learning. First, we consider a special case in which humans can select the better model with 100% accuracy (deterministic selection). Second, we relax this to a probabilistic selection. For the probabilistic selection, we propose an algorithm we call Model Alignment through Probabilistic Bisection (MAPB) and characterize its sample complexity.
@@ 4.1 Deterministic Bisection
The overall goal is to use the least number of samples to find a $\theta$ such that $d(\theta,\theta^{*})\leq\varepsilon$. When a human is able to select a better model with 100% accuracy, the deterministic bisection algorithm can be used to approach the true parameter. The algorithm operates by maintaining a search interval $[\theta^{-},\theta^{+}]$ that contains the true parameter $\theta^{*}$. The interval is initialized by $[-\beta_{\Theta},\beta_{\Theta}]$, and at round $k$, the center of this interval with respect to the distance $d(\cdot,\cdot)$ is used as the query point. Specifically, after defining the query point $\theta_{k}$ to be the single point in the set $H(\theta^{-},\theta^{+})$, we ask the human to compare the utility of two points, denoted by $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$. These two points have equal distance $\Delta/2$ from $\theta_{k}$ in opposite directions, where $\Delta$ is called comparison granularity and is a tuning parameter. For example, for the Euclidean distance, at each round, $\theta_{k}=(\theta^{-}+\theta^{+})/2$; also, $c_{\Delta}^{-}(\theta_{k})=\theta_{k}-\Delta/2$, and $c_{\Delta}^{+}(\theta_{k})=\theta_{k}+\Delta/2$. If the human selects $c_{\Delta}^{-}(\theta_{k})$, then we eliminate the upper half of the search interval by updating $\theta^{+}$ to be equal to $\theta_{k}$; otherwise, we eliminate the bottom half of the search interval by setting $\theta^{-}$ to be equal to $\theta_{k}$. After this bisection step, we continue the process by selecting $\theta_{k+1}$ to be $H(\theta^{-},\theta^{+})$ again, and repeat the above step until the length of the search interval falls below the precision $\varepsilon$. The pseudo-code is presented in Algorithm 1.
@@
Figure 1: An illustrative example. Among two choices $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$, $c_{\Delta}^{-}(\theta_{k})$ would be selected because it is closer to the true parameter. In this case, the interval $(\theta_{k},\theta^{+}]$ is eliminated.
@@
Suppose the human makes the selection with 100% accuracy $(\gamma\to 0)$. Fix $\varepsilon>0$. After $O(\log_{2}\left(\frac{\beta_{\Theta}}{\varepsilon}\right))$ rounds, Algorithm 1 reaches a point $\theta_{k}$ such that $d(\theta_{k},\theta^{*})\leq\varepsilon$.
Note that the comparison granularity has no impact on Algorithm 1 and its computational complexity because the deterministic model assumes 100% accuracy for any $\Delta$. However, this deterministic model of human selection as described exhibits two significant limitations: First, in practical scenarios, humans may not always be able to identify the correct choice, implying that $\gamma>0$; second, the accuracy of human selections is influenced by the comparison granularity $\Delta$. To address these limitations, we introduce a more general probabilistic model in the following section.
@@ 4.2 Probabilistic Bisection
First, we enrich the model to capture the probabilistic nature of human comparisons by noting that, in the RUM, when $\gamma$ is strictly positive, the human’s selection is random, with a certain distribution that favors the correct answer, parameterized by $\gamma$. Precisely, as shown in Lemma 3.7, the human selects $c_{\Delta}^{-}(\theta)$ with probability
$p^{-}_{\gamma}(\theta):=\left[1+\exp\left\{\frac{u\Big{(}c_{\Delta}^{+}(\theta)\Big{)}-u\Big{(}c_{\Delta}^{-}(\theta)\Big{)}}{\gamma}\right\}\right]^{-1}\,,$
and the human selects $c_{\Delta}^{+}(\theta)$ with probability $p^{+}_{\gamma}(\theta)$, which is equal to $1-p^{-}_{\gamma}(\theta)$.
We note that, in this equation, the selection probability depends on the comparison granularity. This dependence naturally captures the difficulty of the selection for the human. On the one hand, when $\Delta$ is too small, $c_{\Delta}^{-}(\theta)$ and $c_{\Delta}^{+}(\theta)$ converge toward each other; hence, $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ converge to $1/2$, representing the difficulty of distinguishing between two highly similar choices. On the other hand, when $\Delta$ is too large, the selection can become difficult as well because both options are far away from $\theta^{*}$. To illustrate, this difficulty can be captured when the distance function converges to a finite value for far away points from $\theta^{*}$, as in Example 3.5. In this case, $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ could also converge to 1/2 (see Example 4.2). As a result of this phenomenon, and for brevity, we exclude $\Delta$ in the notations $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ in the remainder of the paper.
@@
Consider
$u(\theta)=-\frac{1}{1+\exp(-|\theta|)}\,.$
Here, the highest utility is achieved at $\theta=0$. For any finite $\theta$, when $\Delta$ is too large, both $u(c_{\Delta}^{+}(\theta))$ and $u(c_{\Delta}^{-}(\theta))$ are close to -1, and thus $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ are both close to 1/2.
@@
Let $\varepsilon>0$ denote the precision level and let $\delta>0$ represent the confidence level. The $(\varepsilon,\delta)$-alignment problem aims to identify a $\theta$ such that $\mathbb{P}(d(\theta,\theta^{*})\leq\varepsilon)\geq 1-\delta$.
@@
Our objective is to solve the $(\varepsilon,\delta)$-alignment problem with small sample complexity. By selecting appropriate values for these parameters, we can also control other metrics such as risk. The algorithm design is inspired by the probabilistic bisection developed by Frazier et al. (2019). Although their work primarily focuses on the analysis of convergence rates, our emphasis is on conducting a sample complexity analysis for an algorithm designed to terminate when both the precision level and confidence level criteria are met.
@@
If $\theta<\theta^{*}$, then $p_{\gamma}^{-}(\theta)<1/2$ and $p_{\gamma}^{+}(\theta)>1/2$; and vice versa. We introduce several notations to enhance simplicity. Let $Y(\theta)$ denote the selection: $Y(\theta)=c_{\Delta}^{-}(\theta)$ with probability $p_{\gamma}^{-}(\theta)$ and $Y(\theta)=c_{\Delta}^{+}(\theta)$ with probability $p_{\gamma}^{+}(\theta)$. Define $\tilde{Z}(\theta)=2\cdot\boldsymbol{1}(Y(\theta)=c_{\Delta}^{+}(\theta))-1$, so that $\tilde{Z}(\theta)=1$ if $Y(\theta)=c_{\Delta}^{+}(\theta)$ (when we believe $\theta^{*}$ is more likely to be to the right of $\theta$) and $\tilde{Z}(\theta)=-1$ if $Y(\theta)=c_{\Delta}^{-}(\theta)$ (when we believe $\theta^{*}$ is more likely to be to the left of $\theta$). Note that we have $\tilde{Z}(\theta)=1$ with probability $p_{\gamma}^{+}(\theta)$. Also, define
@@
$\tilde{p}(\theta)=\mathbb{P}\big{(}\tilde{Z}(\theta)\text{ correctly identifies the direction}\big{)}=\begin{cases}\mathbb{P}(\tilde{Z}(\theta)=1),&\mbox{if }\theta\leq\theta^{*}\\ \mathbb{P}(\tilde{Z}(\theta)=-1),&\mbox{if }\theta>\theta^{*}\end{cases}\,.$
@@
Figure 2 shows the structure of introducing our main algorithm. Algorithm MAPB, described as follows, involves two types of moves: the horizontal moves and the vertical moves. For any fixed $\theta$, we ask the human to repeatedly evaluate $c_{\Delta}^{-}(\theta)$ and $c_{\Delta}^{+}(\theta)$. We refer to this as the vertical moves. After the vertical stopping criteria are satisfied, the algorithm decides the next $\theta$ to evaluate according to a certain distribution that will be defined. We call this step the horizontal move. The horizontal moves continue until the horizontal stopping criteria are satisfied. In the following two sections, we discuss the vertical moves and horizontal moves, as well as their stopping criteria, in sequence. Then, as shown in Figure 2, we introduce the two-stage Human-AI collaboration framework for multiple dimensions in Section 5 and practical implementations in Section 6.
@@
One-dimensional MAPB (Section 4.2) Horizontal moves (Section 4.2.2); Complexity characterization (Section 4.3) Sample complexity (Theorem 4.5) Extension to multi-dimensional setting (Section 5); Two-stage framework for human-AI collaboration (Section 5.1); Complexity characterization (Section 5.3) Practical implementations (Section 6) Vertical moves (Section 4.2.1); Complexity characterization (Section 4.4)
Figure 2: Roadmap for introducing Algorithm 2.
@@ 4.2.1 Vertical moves
Fix some $\theta$. We repeatedly ask for human evaluations of the query point $\theta$, achieved through comparisons of $c^{+}_{\Delta}(\theta)$ and $c^{-}_{\Delta}(\theta)$. We observe a sequence of signs $\{\tilde{Z}_{i}(\theta)\}_{i}$. If $\theta<\theta^{*}$, then the expectation $\mathbb{E}[\tilde{Z}_{i}(\theta)]=2\tilde{p}(\theta)-1>0$; if $\theta>\theta^{*}$, then the expectation $\mathbb{E}[\tilde{Z}_{i}(\theta)]=1-2\tilde{p}(\theta)<0$. From the response $\tilde{Z}_{i}(\theta)$, we define the simple random walk $\{S_{m}(\theta):m\geq 0\}$ with $S_{0}(\theta)=0$ and $S_{m}(\theta)=\sum_{i=1}^{m}\tilde{Z}_{i}(\theta)$ for $m\geq 1$. Sequential test of power one indicates whether the drift $\omega$ of a random walk satisfies the hypothesis $\omega<0$ versus $\omega>0$.
Define the stopping time $\tau^{\uparrow}_{1}(\theta)=\inf\{m\in\mathbb{N}:|S_{m}(\theta)|\geq\hbar_{m}\}$, where $\{\hbar_{m}\}_{m}$ is a positive sequence. Here, $\uparrow$” and $\rightarrow$” represent vertical and horizontal, respectively. The test decides that $\theta<\theta^{*}$, if $S_{\tau_{1}^{\uparrow}(\theta)}(\theta)\geq\hbar_{\tau^{\uparrow}_{1}(\theta)}$; it decides that $\theta>\theta^{*}$ if $S_{\tau_{1}^{\uparrow}(\theta)}(\theta)\leq-\hbar_{\tau^{\uparrow}_{1}(\theta)}$; and the test does not make a decision if $\tau^{\uparrow}_{1}(\theta)=\infty$.
Assume now that the algorithm queries a point $\theta_{k}$ (not equal to $\theta^{*}$) at the $(k+1)^{th}$ iteration. We then observe a random walk with the $m^{th}$ term $S_{k,m}=S_{k,m}(\theta_{k})=\sum_{i=1}^{m}Z_{k,i}(\theta_{k})$ until we reach the end of the test, which is called the power-one test (Robbins and Siegmund 1974). Define the new signal:
@@
Define $\hbar_{m}=(2m(\ln(m+1)-\ln\eta))^{1/2}$. In the event that $\theta_{k}>\theta^{*}$, $\mathbb{P}(Z_{k}(\theta_{k})=+1|\theta_{k})\leq\eta/2$. In the event that $\theta_{k}<\theta^{*}$, $\mathbb{P}(Z_{k}(\theta_{k})=-1|\theta_{k})\leq\eta/2$. That is, the detection accuracy $p(\theta)$ when the vertical move stops is bounded by $p(\theta)\geq 1-\eta/2$.
Lemma 4.4 states that the detection accuracy is lower bounded by $p_{c}:=1-\eta/2$, where the accuracy can be chosen through the selection of $\eta$. As $\theta$ approaches $\theta^{*}$, differentiating between choices becomes increasingly challenging because the drift tends to diminish, potentially requiring more steps in the testing process. However, our objective, you must recall, is to locate a $\theta$ such that $|\theta-\theta^{*}|\leq\varepsilon$, with a confidence level of $1-\delta$. In the vicinity of $\theta^{*}$, waiting until $|S_{m}(\theta)|\geq\hbar_{m}$ to terminate the test may not be necessary. A high volume of queries at a particular point $\theta_{k}$ suggests close proximity between $\theta_{k}$ and $\theta$, so that considering a earlier termination of the test may be advisable.
Based on this insight, we introduce a stopping time $\tau^{\uparrow}(\delta,\varepsilon)$, which is determined by the precision and confidence levels. This stopping time regulates the maximum number of queries carried out at a single point as $\theta$ approaches the local neighborhood of the true parameter (a concept we elaborate on in Section 4.4). Specifically, $\theta$ converges toward $\theta^{*}$ in a defined local neighborhood, denoted as $\{\theta^{\prime}:d(\theta,\theta^{\prime})\leq a\}$ for some positive constant $a$; here, surpassing the threshold $\tau^{\uparrow}(\delta,\varepsilon)$ signifies the difficulty in distinguishing between the parameters because of their proximity, and the algorithm terminates.